### v1

# Communication-Efficient Distributed Deep Learning via Federated Dynamic Averaging

Abstract.
Driven by the ever-growing volume and decentralized nature of data, coupled with the escalating size of modern models, distributed deep learning (DDL) has been entrenched as the preferred paradigm for training. However, frequent synchronization of DL models, encompassing millions to many billions of parameters, creates a communication bottleneck, severely hindering scalability. Worse yet, DDL algorithms typically waste valuable bandwidth, and make themselves less practical in bandwidth-constrained federated settings, by relying on overly simplistic, periodic, and rigid synchronization schedules. To address these shortcomings, we propose Federated Dynamic Averaging (FDA), a communication-efficient DDL strategy that dynamically triggers synchronization based on the value of the model variance. Through extensive experiments across a wide range of learning tasks we demonstrate that FDA reduces communication cost by orders of magnitude, compared to both traditional and cutting-edge communication-efficient algorithms. Remarkably, FDA achieves this without sacrificing convergence speed — in stark contrast to the trade-offs encountered in the field. Additionally, we show that FDA maintains robust performance across diverse data heterogeneity settings.
1. Introduction
Deep Learning (DL) stands at the forefront of the most significant technological advancements in recent times, with Deep Neural Networks (DNNs) achieving remarkable success in fields such as computer vision (Kolesnikov et al., 2020), natural language processing (Devlin et al., 2019; Brown et al., 2020) and speech recognition (Amodei et al., 2015). These rapid advancements coincide with the big data era, marked by an unprecedented scale of training datasets (Sun et al., 2017; Russakovsky et al., 2015). These datasets are not only growing in size but are often physically distributed and cannot be easily centralized due to business considerations, privacy concerns, bandwidth limitations, and data sovereignty laws (Kairouz et al., 2019; McMahan et al., 2016; Konečný et al., 2016). In response, Distributed Deep Learning (DDL) has emerged as an alternative paradigm to the traditional centralized approach (Zinkevich et al., 2010; Peteiro-Barral and Guijarro-Berdiñas, 2012; Chilimbi et al., 2014), offering efficient handling of large-scale data across multiple worker-nodes, enhancing the speed of training DL models and paving the way for more scalable and resilient DL applications (Moritz et al., 2016; Abadi et al., 2015; Li et al., 2020).
Most DDL methods are iterative, where, in each iteration, some amount of local training is followed by synchronization of the local model with the global one. The predominant method, based on the bulk synchronous parallel (BSP) approach (Valiant, 1990), is to average the local model updates and then apply the average update to each local model (Zinkevich et al., 2010; Dekel et al., 2010). Less synchronized variants have also been proposed, to ameliorate the effect of straggler workers (Niu et al., 2011; Cui et al., 2014; Gupta et al., 2016; Mitliagkas et al., 2016; Dutta et al., 2018; Ho et al., 2013) but sometimes compromise convergence speed and model quality.
A critical challenge inherent in the traditional techniques, especially in federated DL settings, where models are huge and worker interconnections are slow, is the communication bottleneck, restricting system scalability (Shi and Chu, 2017; Tang et al., 2023). This leads to a low computation-to-communication ratio (Shi et al., 2021; Lin et al., 2017). Addressing the communication overhead to expedite DDL algorithms has been a focal point of research for many years; speeding-up SGD is arguably among the most impactful and transformative problems in machine learning (Wang and Joshi, 2021).
The most direct method to alleviate the communication burden is to reduce the frequency of communication rounds. Local-SGD is the prime example of this approach. It allows workers to perform $\tau$ local update steps on their models before aggregating them, as opposed to averaging the updates in every step (Zhang et al., 2016; Yu et al., 2018; Haddadpour et al., 2019). Although Local-SGD is effective in reducing communication while maintaining similar accuracy levels (Smith et al., 2016; Wang and Joshi, 2021), determining the optimal value of $\tau$ presents a significant challenge, with only a handful of studies offering theoretical insights into its influence on convergence (Wang and Joshi, 2021; Stich, 2019; Zhou and Cong, 2017; Yu et al., 2018).
To further reduce communication costs of Local-SGD, more sophisticated communication strategies introduce varying sequences of local update steps $\{\tau_{0},...,\tau_{R}\}$, instead of a fixed $\tau$. In (Wang and Joshi, 2018), in order to minimize convergence error with respect to wall-time, the authors proposed a decreasing sequence of local update steps. Conversely, the focus in (Haddadpour et al., 2019) was on reducing the number of communication rounds for a fixed number of model updates and prescribed an increasing sequence. These contrasting approaches underscore the multifaceted nature of communication strategies in distributed deep learning, highlighting not only the absence of a one-size-fits-all solution but also the growing need for dynamic, context-aware strategies that can continuously adapt to the specific intricacies of the learning task.
Main Contributions. Our work addresses critical efficiency challenges in DDL, particularly in communication-constraint environments, such as the ones encountered in Federated Learning (FL) (Kairouz et al., 2019) applications. We introduce Federated Dynamic Averaging (FDA), a novel, adaptive distributed deep learning strategy, massively improving communication efficiency over previous work. Our contributions are as follows:
  • FDA dynamically decides to synchronize local nodes when model variance across nodes exceeds a threshold. This strategy drastically reduces communication requirements, while preserving cohesive progress towards the shared training objective.
  • We evaluate and compare FDA with other DDL algorithms through a comprehensive suite of experiments with diverse datasets, models, and tasks. This extensive approach—requiring over $200$K GPU hours—compensates for the stochasticity associated with training DNNs, thereby enabling us to draw more definitive conclusions.
  • We demonstrate that FDA outperforms traditional and contemporary FL algorithms by orders of magnitude in communication, while maintaining equivalent model performance. Furthermore, it effectively balances the competing demands of reducing communication and computation, providing greatly improved trade-offs.
  • We show that FDA remains robust and efficient under various data distributions (Non-IID settings), maintaining comparable performance to the IID case. Furthermore, we demonstrate that FDA leads to models with excellent generalization, inherently mitigating overfitting exhibited by previous techniques.
2. Related Work
Problem formulation. Consider distributed training of deep neural networks over multiple workers (Dean et al., 2012; Li et al., 2021). In this setting, each worker has access to its own set of training data $\mathcal{D}_{k}$, and the collective goal is to find a common model $\mathbf{w}\in\mathbb{R}^{d}$ by minimizing the overall training loss. This scenario can be effectively modeled as a distributed optimization problem, formulated as follows:
(1) $\underset{\mathbf{w}\in\mathbb{R}^{d}}{\mathrm{minimize}}\;\;F(\mathbf{w})\triangleq\frac{1}{K}\sum_{k=1}^{K}F_{k}(\mathbf{w})$
where $K$ is the number of workers and $F_{k}(\mathbf{w})\triangleq\mathbb{E}_{\zeta_{k}\sim\mathcal{D}_{k}}\left[\ell(\mathbf{w};\zeta_{k})\right]$ is the local objective function for worker $k$. Function $\ell(\mathbf{w};\zeta_{k})$ represents the loss for data sample $\zeta_{k}$ given model $w$.
Communication efficient Local-SGD. The work in (Li et al., 2021) decomposes each round into two phases. In the first phase, each worker runs Local-SGD with $\tau=I_{1}$, while the second phase run $I_{2}$ steps with $\tau=1$(Li et al., 2021) proposes to exponentially decay $I_{1}$ every $M$ rounds. In the heterogeneous setting, the work in (Qin et al., 2023), by analysing the convergence rate, proposes an increasing sequence of local update steps for strongly-convex local objectives, and suggests fixed local steps for convex, non-convex local objectives. The study in (Yu and Jin, 2019) dynamically increases batch sizes to reduce communication rounds, maintaining the same convergence rate as SSP-SGD. However, the large-batch approach leads to poor generalization (Lin et al., 2018; Hoffer et al., 2018; Shallue et al., 2018), a challenge addressed by the post-local SGD method (Lin et al., 2018), which divides training into two phases: BSP-SGD followed by Local-SGD with a fixed number of steps. In the Lazily Aggregated Algorithm (LAG) (Chen et al., 2018), a different approach was taken, using only new gradients from some selected workers and reusing the outdated gradients from the rest, which essentially skips communication rounds.
Federated Averaging (FedAvg) (McMahan et al., 2016) is another representative communication efficient Local-SGD algorithm, which is a pivotal method in Federated Learning (FL) (Kairouz et al., 2019). In the FL setting with edge computing systems, the work in (Wang et al., 2019) tries to find the optimal synchronization period $\tau$ subject to local computation and aggregation constraints. Recently (Mills et al., 2023), in the FL setting with the assumption of strongly-convex objectives, by analysing the balance between fast convergence and higher-round completion rate, a decaying local update step scheme emerged.
Accelerating convergence. An indirect, yet highly effective, way to mitigate the communication burden in DDL, is to speed up convergence. Consequently, recent works have deployed accelerated versions of SGD to the distributed setting. Specifically, FedAdam (Reddi et al., 2021) extends Adam (Kingma and Ba, 2017) and FedAvgM (Hsu et al., 2019) extends SGD with momentum (SGD-M) (Sutskever et al., 2013a). Recently, Mime (Karimireddy et al., 2021) provides a framework to adapt arbitrary centralized optimization algorithms to the FL setting. However, these methods still suffer from the model divergence problem, particularly in heterogeneous settings. When solving (1), the disparity between each worker’s optimal solution $\mathbf{w}_{k}^{*}$ for their objective $F_{k}$, and the global optimum $\mathbf{w}^{*}$ for $F$, can potentially cause worker models to diverge (drift) towards their disparate minima (Karimireddy et al., 2019; Reddi et al., 2021). The result is slow and unstable convergence with significant communication overhead. To address this problem, the SCAFFOLD algorithm (Karimireddy et al., 2019) used control-variates (in the same spirit to SVRG (Johnson and Zhang, 2013)), with significant speed-up. FedProx (Sahu et al., 2018) re-parametarized FedAvg (McMahan et al., 2016) by adding $L^{2}$ regularization in the workers’ objectives to be near the global model. Lastly, FedDyn (Acar et al., 2021) improved upon these ideas with a dynamic regularizer making sure that if local models converge to a consensus, this consensus point aligns with the stationary point of the global objective function.
Compression. To reduce communication overhead in DDL, significant efforts have been directed towards minimizing message sizes. Key strategies include sparsification, where only crucial components of information are transmitted, as explored in the works (Ström, 2015; Aji and Heafield, 2017; Alistarh et al., 2018), and quantization techniques, which involve transmitting only quantized gradients, as detailed in studies like (Rabbat and Nowak, 2005; Shlezinger et al., 2020). Notably, these compression techniques can be effectively combined with Local-SGD methods to enhance communication-efficiency further. A prime example is Qsparse-local-SGD (Basu et al., 2019), which integrates aggressive sparsification and quantization with Local-SGD, achieving substantial communication savings to reach specific target accuracies. A comprehensive survey of compression techniques in DDL can be found in (Wang et al., 2023).
3. Federated Dynamic Averaging
Notation. At each time instance $t$, each worker $k$ independently maintains its own vector of model parameters, [^1] The terms “model” and ”model parameters” are used interchangeably, as is common in the literature. denoted as $\mathbf{w}_{t}^{(k)}\in\mathbb{R}^{d}$. Let $\mathbf{\Delta w}^{(k)}_{t}$ be the update vector (Goodfellow et al., 2016) computed by some optimization algorithm (e.g., SGD, Adam) encompassing the learning rate and relevant gradients. In a distributed learning step $t$, each worker $k$ applies the following update:
(2) $\mathbf{w}^{(k)}_{t+1}=\mathbf{w}^{(k)}_{t}+\mathbf{\Delta w}^{(k)}_{t}$
Let $\overline{\mathbf{w}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{w}_{t}^{(k)}$ represent the average model at time $t$ and $\overline{\mathbf{w}}_{t_{0}}$ the initial model shared among the workers at the start of the current communication round. Moreover, we introduce the drift vector $\mathbf{u}_{t}^{(k)}=\mathbf{w}_{t}^{(k)}-\overline{\mathbf{w}}_{t_{0}}$, that is, the change to the local model of the $k$-th worker at time $t$ since the beginning of the current round at time $t_{0}$, and $\overline{\mathbf{u}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{u}_{t}^{(k)}=\overline{\mathbf{w}}_{t}-\overline{\mathbf{w}}_{t_{0}}$ as the average drift.
Model Variance. The model variance quantifies the dispersion or spread of the worker models around the average model. At time $t$, it is defined as:
$\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$ $=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{w}_{t}^{(k)}-\overline{\mathbf{w}}_{t}\right\|_{2}^{2}=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{w}_{t}^{(k)}-\mathbf{w}_{t_{0}}+\mathbf{w}_{t_{0}}-\overline{\mathbf{w}}_{t}\right\|_{2}^{2}$
(3) $=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}^{(k)}_{t}-\overline{\mathbf{u}}_{t}\right\|_{2}^{2}=...=\left(\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}^{(k)}_{t}\right\|_{2}^{2}\right)-\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$
This measure provides insight into how closely aligned the workers’ models are at any given time. High variance indicates that the models are widely spread out, essentially drifting apart, leading to a lack of cohesion in the aggregated model. Conversely, a moderate-low variance suggests that the workers’ models are closely aligned, working collectively towards the shared objective. The variance plays a critical role in our approach, as it helps us gauge the state of the training process making informed decisions about round termination.
Round Invariant. We introduce the Round Invariant (RI), a condition that bounds model variance below a threshold $\Theta$:
(4) $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\leq\Theta$
Our algorithm continuously monitors the model variance to maintain the RI. If the variance estimate exceeds the threshold $\Theta$ at any point $t$, the training round immediately terminates. This triggers synchronization, where all local model weights $\mathbf{w}^{(k)}_{t}$ are averaged and consolidated into the new global model $\overline{\mathbf{w}}_{t}$.
Monitoring the RI. At this point, our focus shifts to devising a method to monitor the RI, as defined in (4). We begin by defining the local state for each worker $k$, denoted by $S_{k}(t)\in\mathbb{R}^{p}$, which encapsulates information from worker $k$ at time $t$; it is updated arbitrarily. Next, we introduce the global state, $S(t)\in\mathbb{R}^{p}$, which represents the collective state of the distributed system at time $t$.
(5) $S(t)=\frac{1}{K}\sum_{k=1}^{K}S_{k}(t)$
Lastly, we define a special non-linear function $H:\mathbb{R}^{p}\to\mathbb{R}$, with the following property:
(6) $\;H(S(t))\leq\Theta\;\Rightarrow\;\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\leq\Theta\quad\forall\;t$
Conceptually, if we define $S_{k}(t)=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\mathbf{u}_{t}^{(k)}\right)\in\mathbb{R}\times\mathbb{R}^{d}$ and $H(v,x)=v-\left\|x\right\|_{2}^{2}$, then, it immediately follows from (3), (5) that $H\left(S\left(t\right)\right)=\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$ and property (6) is satisfied. Direct monitoring of the RI is hindered by the high dimensionality of $S_{k}(t)$, with $d$ potentially ranging from millions to many billions, values commonly found in DNNs. To mitigate this communication burden, we must apply dimensionality reduction to the local drifts $\mathbf{u}_{t}^{(k)}$ and identify an appropriate function $H$, accordingly. However, this reduction in information within local states causes RI monitoring to become approximate. Upcoming sections will detail two strategies–termed ”linear”, and ”sketch”–each offering a different balance between communication efficiency and approximation accuracy.
Algorithm. The proposed FDA algorithm is formalized in Algorithm 1. To streamline notation, $t$ also denotes each training round, which unfolds into three main parts:
  1. (1)
    Broadcast: Model $\mathbf{w}_{t}$ is distributed to all workers (Line 3).
  2. (2)
    Local Training: As long as we can guarantee the RI, workers continue training concurrently (Lines 4-9). Specifically, workers perform in-parallel updates to their local models $\mathbf{w}^{(k)}_{t}$ using mini-batch SGD, or some other optimization algorithm, by processing their local data (Lines 5-7). After each update, workers compute and broadcast their local states $S_{k}(t)$ (Line 8). Then, the global state $S(t)$ is calculated via (5), and using the function $H$, an estimate of the model variance is derived and compared with the threshold $\Theta$ (Line 9). Depending on this comparison, the local training stage either repeats or terminates.
  3. (3)
    Model Averaging: This stage is triggered when we can no longer guarantee the RI, i.e., $H(S(t))>\Theta$. At this point, all local models are communicated and then averaged to form $\mathbf{w}_{t+1}$, that is, the next round’s global model (Lines 10-12).
Algorithm 1 Federated Dynamic Averaging - FDA
1:Require: $S_{k}(t)$: The local state $S_{k}(t)\in\mathbb{R}\times\mathbb{R}^{p}$ with $p\ll d$
2:Require: $H(x)$: A function s.t. $H(S(t))\leq\Theta\Rightarrow\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\leq\Theta$
3:Require: $K$: The number of workers indexed by $k$
4:Require: $\Theta$: The model variance threshold
5:Require: $b$: The local mini-batch size
6:
7:Initialize $\mathbf{w}_{1}\in\mathbb{R}^{d}$
8:for each round $t=1,2,\dots$ do
9:     communicate $\mathbf{w}_{t}$ to all workers
10:     repeat
11:          for each worker $k=1,\dots,K$ in parallel do
12:               $\mathcal{B}_{k}\leftarrow$ (sample a batch of size $b$ from $\mathcal{D}_{k}$)
13:               $\mathbf{w}^{(k)}_{t}\leftarrow\mathbf{w}^{(k)}_{t}-\eta\nabla\ell(\mathbf{w}^{(k)}_{t};\mathcal{B}_{k})$triangleright In-place
14:          communicate $S_{k}(t)$
15:     until $H(S(t))>\Theta$triangleright As per (5)
16:     for each worker $k=1,\dots,K$ in parallel do
17:          communicate $\mathbf{w}^{(k)}_{t}$
18:     $\mathbf{w}_{t+1}\leftarrow\frac{1}{K}\sum_{k=1}^{K}\mathbf{w}^{(k)}_{t}$
19:
3.1. Linear Approximation
[…]
$S_{k}(t)=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\right)\in\mathbb{R}\times\mathbb{R}\;\;,\;\;\left\|\xi\right\|_{2}=1$
$H\left(v\,,\,x\right)=v-x^{2}$
Then, the condition $H\left(S\left(t\right)\right)\leq\Theta$ implies $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\leq\Theta$, since:
$H\left(S\left(t\right)\right)$ =(5)${=}H\left(\frac{1}{K}\sum_{k=1}^{K}S_{k}(t)\right)=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left|\frac{1}{K}\sum_{k=1}^{K}\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\right|^{2}$
$=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left|\left\langle\xi\,,\,\overline{\mathbf{u}}_{t}\right\rangle\right|^{2}\geq\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}=\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$
Furthermore, a random choice of $\xi$ is likely to perform poorly (terminate a round prematurely), as it is likely to be close to orthogonal to $\overline{\mathbf{u}}_{t}$. A good choice would be a vector correlated to $\overline{\mathbf{u}}_{t}$. A heuristic choice is to take $\overline{\mathbf{u}}_{t_{0}}$, i.e., the change vector right before the current round started. All workers can estimate this without communication cost as the difference of the models at the beginning of the current and previous rounds and use $\xi=\frac{w_{t_{0}}-w_{t_{-1}}}{\left\|w_{t_{0}}-w_{t_{-1}}\right\|_{2}}.$
3.2. Sketch Approximation
A better estimator for $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$ can be obtained through the utilization of AMS sketches, as detailed in (Cormode and Garofalakis, 2005). An AMS sketch of a vector $\mathbf{v}\in\mathbb{R}^{d}$ is an $l\times m$ real matrix $\Xi$:
$\operatorname{sk}\left(\mathbf{v}\right)=\Xi=\begin{bmatrix}\xi_{1}&\xi_{2}&\dots&\xi_{l}\end{bmatrix}^{\top}\in\mathbb{R}^{l\times m}\;\;,\;\;l\cdot m\ll d$
The $\operatorname{sk}(\cdot)$ operator is linear and can be computed in $\mathcal{O}\left(l\cdot d\right)$ steps. Let $l=\mathcal{O}\left(\log 1/\delta\right)$ and $m=\mathcal{O}\left(1/\epsilon^{2}\right)$. The function
$\mathcal{M}_{2}\left(\Xi\right)=\mathrm{median}\left\{\left\|\xi_{i}\right\|_{2}^{2}\;,\;i=1,\dots,l\right\}$
is an excellent estimator of $\left\|\mathbf{v}\right\|_{2}^{2}$, as with probability at least $\left(1-\delta\right)$, $\mathcal{M}_{2}\left(\operatorname{sk}\left(\mathbf{v}\right)\right)$ is within $\epsilon$-relative error of $\left\|\mathbf{v}\right\|_{2}^{2}$. We then define:
$S_{k}(t)=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\right)\in\mathbb{R}\times\mathbb{R}^{l\times m}$
$H\left(\mathbf{v}\,,\,\Xi\right)=\mathbf{v}-\frac{1}{1+\epsilon}\mathcal{M}_{2}\left(\Xi\right)$
Then, the condition $H\left(S\left(t\right)\right)\leq\Theta$ implies $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\leq\Theta$ with probability at least $\left(1-\delta\right)$. To prove this, we proceed as follows: we first observe $\mathcal{M}_{2}\left(\frac{1}{K}\sum_{k=1}^{K}\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\right)=\mathcal{M}_{2}\left(\operatorname{sk}\left(\frac{1}{K}\sum_{k=1}^{K}\mathbf{u}_{t}^{(k)}\right)\right)=\mathcal{M}_{2}\left(\operatorname{sk}\left(\overline{\mathbf{u}}_{t}\right)\right)\leq(1+\epsilon)\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$ with probability at least $(1-\delta)$. Then:
$H\left(S\left(t\right)\right)$ $=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\frac{1}{1+\epsilon}\mathcal{M}_{2}\left(\frac{1}{K}\sum_{k=1}^{K}\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\right)$
$\geq\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}=\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\;\text{with prob. at least}\;(1-\delta)$
4. Experiments
4.1. Setup
Platform. We employ TensorFlow (Abadi et al., 2015), integrated with Keras (Chollet et al., 2015), as the platform for conducting our experiments. We modified TensorFlow to include our FDA variants and all competitive algorithms.
Table 1. Summary of Experiments
Hyper-Parameters Training NN d Dataset $\Theta$ b K Optimizer Algorithms LeNet-5 62K MNIST $\{0.5,1,1.5,2,3,5,7\}$ 32 { 5, 10, …, 60 } Adam FDA, Synchronous, FedAdam VGG16* 2.6M MNIST $\{20,25,30,50,75,90,100\}$ 32 { 5, 10, …, 60 } Adam FDA, Synchronous, FedAdam DenseNet121 6.9M CIFAR-10 $\{200,250,275,300,325,350,400\}$ 32 { 5, 10, …, 30 } SGD-NM FDA, Synchronous, FedAvgM DenseNet201 18M CIFAR-10 $\{350,500,600,700,800,850,900\}$ 32 { 5, 10, …, 30 } SGD-NM FDA, Synchronous, FedAvgM (fine-tuning) ConvNeXtLarge 198M CIFAR-100 $\{25,50,100,150\}$ 32 { 3, 5 } AdamW FDA, Synchronous
Datasets & Models. The core experiments involve training Convolutional Neural Networks (CNNs) of varying sizes and complexities on two datasets: MNIST (Deng, 2012) and CIFAR-10 (Krizhevsky, 2012). For the MNIST dataset, we employ LeNet-5 (Lecun et al., 1998), composed of approximately 62 thousand parameters, and a modified version of VGG16 (Simonyan and Zisserman, 2014), denoted as VGG16*, consisting of 2.6 million parameters. VGG16* was specifically adapted for the MNIST dataset, a less demanding learning problem compared to ImageNet (Russakovsky et al., 2015), for which VGG16 was designed. In VGG16*, we omitted the $512$-channel convolutional blocks and downscaled the final two fully connected (FC) layers from $4096$ to $512$ units each. Both models use Glorot uniform initialization (Glorot and Bengio, 2010). For CIFAR-10, we utilize DenseNet121 and DenseNet201 (Huang et al., 2016), as implemented in Keras (Chollet et al., 2015), with the addition of dropout regularization layers at rate 0.2 and weight decay of $10^{-4}$, as prescribed in (Huang et al., 2016). The DenseNet121 and DenseNet201 models have 6.9 million and 18 million parameters, respectively, and are both initialized with He normal (He et al., 2015).
Lastly, we explore a transfer learning scenario on the CIFAR-100 dataset (Krizhevsky, 2012), a choice reflecting the DL community’s growing preference of using pre-trained models in such downstream tasks (Han et al., 2021; Ridnik et al., 2021). For example, a pre-trained visual transformer (ViT) on ImageNet, transferred to classify CIFAR-100, is currently on par with the state-of-the-art results for this task (Dosovitskiy et al., 2020). We adopt this exact transfer learning scenario, leveraging the more powerful ConvNeXtLarge model, pre-trained on ImageNet, with $198$ million parameters (Liu et al., 2022; Chollet et al., 2015). Following the feature extraction step (Goodfellow et al., 2016; Zhuang et al., 2019), the testing accuracy on CIFAR-100 stands at $60\%$. Subsequently, we employ and evaluate our FDA algorithms in the arduous fine-tuning stage, where the entirety of the model is trained (Yosinski et al., 2014).
Algorithms. We consider five distributed deep learning algorithms: LinearFDA, SketchFDA, Synchronous  [^2] The name was derived from the Bulk Synchronous Parallel approach; can be understood as a special case of the FDA Algorithm 1 where $\Theta$ is set to zero., FedAdam (Reddi et al., 2021), and FedAvgM (Hsu et al., 2019); the first three are standard in all experiments. Depending on the local optimizer, Adam (Kingma and Ba, 2017) or SGD with Nesterov momentum (SGD-NM) (Sutskever et al., 2013b), we also include their federated counterparts FedAdam or FedAvgM, respectively.
Evaluation Methodology. Comparing DDL algorithms is not straightforward. For example, comparing DDL algorithms based on the average cost of a training epoch can be misleading, as it does not consider the effects on the trained model’s quality. To achieve a comprehensive performance assessment of FDA, we define a training run as the process of executing the DDL algorithm under evaluation, on (a) a specific DL model and training dataset, and (b) until a final epoch in which the trained model achieves a specific testing accuracy (termed as Accuracy Target in the Figures). Based on this definition, we focus on two performance metrics:
  1. (1)
    Communication cost, which is the total data (in bytes) transmitted by all workers. Translating this cost to wall-clock time (i.e., the total time required for the computation and communication of the DDL) depends on the network infrastructure connecting the workers. Its impact is larger in FL scenarios, where workers often use slower Wi-Fi connections.
  2. (2)
    Computation cost, which is the number of mini-batch steps (termed as In-Parallel Learning Steps in the Figures) performed by each worker. Translating this cost to wall-clock time is determined by the mini-batch size and the computational resources of the nodes. Its impact is larger for nodes with lower computational resources.
Hyper-Parameters & Optimizers. Hyper-parameters unique to each training dataset and model are detailed in Table 1; $\Theta$ is pertinent to FDA algorithms and not applicable to others. Notably, a guideline for setting the parameter $\Theta$ is provided in Section 4.2. For experiments involving FedAvgM and FedAdam, we use $E=1$ local epochs, following (Reddi et al., 2021). For experiments with LeNet-5 and VGG16*, local optimization employs Adam, using the default settings as per (Kingma and Ba, 2017). In these cases, FedAdam also adheres to the default settings for both local and server optimization (Reddi et al., 2021; Chollet et al., 2015). For DenseNet121 and DenseNet201, local optimization is performed using SGD with Nesterov momentum (SGD-NM), setting the momentum parameter at $0.9$ and learning rate at $0.1$ (Huang et al., 2016). For FedAvgM, local optimization is conducted with default settings (Hsu et al., 2019; Chollet et al., 2015), while server optimization employs SGD with momentum, setting the momentum parameter and learning rate to $0.9$ and $0.316$, respectively (Reddi et al., 2021). Lastly, for the transfer learning experiments, local optimization leverages AdamW (Loshchilov and Hutter, 2017), with the hyper-parameters used for fine-tuning ConvNeXtLarge in the original study (Liu et al., 2022).
Data Distribution. In all experiments, the training dataset is divided into approximately equal parts among the workers. To assess the impact of data heterogeneity, we explore three scenarios:
  1. (1)
    IID — Independent and identically distributed.
  2. (2)
    Non-IID: $X$— A portion $X\%$ of the dataset is sorted by label and sequentially allocated to workers, with the remainder distributed in an IID fashion.
  3. (3)
    Non-IID: Label $Y$ — All samples from label $Y$ are assigned to a few workers, while the rest are distributed in an IID manner.
Main Findings. The main findings of our experimental analyses are: (1) LinearFDA and SketchFDA outperform the Synchronous, FedAdam and FedAvgM techniques (their use depends on the local optimizer choice) by 1-2 orders of magnitude in communication, while maintaining equivalent model performance. (2) LinearFDA and SketchFDA also significantly outperform the FedAdam and FedAvgM techniques in terms of computation. (3) The performance of LinearFDA and SketchFDA is comparable in most experiments. SketchFDA provides a more accurate estimator of the variance and leads to fewer synchronizations than LinearFDA, but has a larger communication overhead for its local state (a sketch, compared to a number). SketchFDA significantly outperforms LinearFDA at the transfer learning scenario. (4) The FDA variants remain robust at various data heterogeneity settings, maintaining comparable performance to the IID case.
4.2. Results
Due to the extensive set of unique experiments (over $1000$), as detailed in Table 1, we leverage Kernel Density Estimation (KDE) plots (Waskom, 2021) to visualize the bivariate distribution of computation and communication costs incurred by each strategy for attaining the Accuracy Target. These KDE plots provide a high-level overview of the cost trade-off for training accurate models. As an illustrative example, Figure 1 depicts the strategies’ bivariate distribution for the LeNet-5 model trained on MNIST with different data heterogeneity setups. In these plots, the SketchFDA distribution is generated from experiments across all hyper-parameter combinations ($\Theta$ and $K$ in Table 1) that attained the Accuracy Target of $0.985$.
Figure 1. LeNet-5 on MNIST
Figure 2. VGG16* on MNIST
Figure 3. DenseNet121 on CIFAR-10
Figure 4. DenseNet201 on CIFAR-10
FDA balances Communication vs. Computation. DDL algorithms face a fundamental challenge: balancing the competing demands of computation and communication. Frequent communication accelerates convergence and potentially improves model performance, but incurs higher network overhead, an overhead that may be prohibitive when workers communicate through lower speed connections. Conversely, reducing communication saves bandwidth but risks hindering, or even stalling, convergence. Traditional DDL approaches, like Synchronous, require synchronizing model parameters after every learning step, leading to significant communication overhead but facilitating faster convergence (lower computation cost). This is evident in Figures 12,  3, and 4 (where Synchronous appears in the bottom right — low computation, very high communication). Conversely, Federated Optimization (FedOpt) methods (Reddi et al., 2021) are designed to be communication-efficient, reducing communication between devices (workers) at the expense of increased local computation. Indeed, as shown in Figures 1-4, FedAvgM and FedAdam reduce communication by orders of magnitude but at the price of a corresponding increase in computation. Our two proposed FDA strategies achieve the best of both worlds: the low computation cost of traditional methods and the communication efficiency of FedOpt approaches, as seen in Figures 1-4. In fact, they significantly outperform FedAvgM and FedAdam in their element, that is, communication-efficiency. Across all experiments, the FDA methods’ distributions lie in the desired bottom left quadrant — low computation, very low communication.
FDA counters diminishing returns. The phenomenon of diminishing returns states that as a DL model nears its learning limits for a given dataset and architecture, each additional increment in accuracy may necessitate a disproportionate increase in training time, tuning, and resources (Goodfellow et al., 2016; Thompson et al., 2021). We first clearly notice this with VGG16* on MNIST in Figure 2 for all three data heterogeneity settings. For a $0.001$ increase in accuracy (effectively 10 misclassified testing images) FedAdam needs an order of magnitude more computation and communication. Similarly, Synchronous requires comparable increases in computation and approximately half an order of magnitude more in communication. On the other hand, the FDA methods suffer a slight (if any) increase in computation and communication for this accuracy enhancement. For DenseNet121 and DenseNet201 on CIFAR-10 (Figures  34), FedAvgM and Synchronous require half an order of magnitude more computation and communication to achieve the final marginal accuracy gains ($0.78$ to $0.81$ for DenseNet121, and $0.78$ to $0.8$ for DenseNet201). In contrast, the FDA methods have almost no increase in communication and comparable increase in computation.
FDA is resilient to data heterogeneity. In DDL, data heterogeneity is a prevalent challenge, reflecting the complexity of real-world applications where the IID assumption often does not hold. The ability of DDL algorithms to maintain consistent performance in the face of non-IID data is a critical metric for their effectiveness and adaptability. Our empirical investigation reveals the FDA methods’ noteworthy resilience in such heterogeneous environments. For LeNet-5 on MNIST, as illustrated in Figure 1, the computation and communication costs required to attain a test accuracy of $0.985$ show negligible differences across the IID and the two Non-IID settings (Label ”00”, $60\%$). Similarly, for VGG16* on MNIST, Figure 2 demonstrates that achieving a test accuracy of $0.995$ incurs comparable computation and communication costs across the IID and the two Non-IID settings (Label ”00”, Label ”$8$”); while overall costs are aligned, the distributions of the computation costs exhibit greater variability, yet remain closely consistent with the IID scenario.
FDA generalizes better. The factors determining how well a DL algorithm performs are its ability to: (1) make the training accuracy high, and (2) make the gap between training and test accuracy small. These two factors correspond to the two central challenges in DL: underfitting and overfitting (Goodfellow et al., 2016). For DenseNet121 on CIFAR-10, with a test accuracy target of $0.8$, as illustrated in Figure 5, Synchronous and FedAvgM exhibit overfitting, with a noticeable discrepancy between training and test accuracy. In stark contrast, the FDA methods have an almost zero accuracy gap. Please note that LinearFDA and SketchFDA reach the test accuracy target of $0.8$ much earlier (at epochs 86 and 91, respectively). Turning our focus to DenseNet201 on CIFAR-10, with a test accuracy target of $0.78$, Synchronous again tends towards overfitting, while FedAvgM shows a slight improvement but still does not match the FDA methods, which continue to exhibit exceptional generalization capabilities, evidenced by a minimal training-test accuracy gap, as shown at Figure 5. Notably, given the necessity to fix hyper-parameters $\Theta$ and $K$ for the training accuracy plots, we selected two representative examples. The patterns of performance we highlighted are consistent across most of the conducted tests.
Figure 5. Training accuracy progression with a test accuracy target (horizontal line) of $0.8$ (left), and $0.78$ (right). Dashed and doted lines indicate when LinearFDA and SketchFDA attain the target accuracy, respectively. A smaller final gap between training and target accuracy indicates less overfitting, i.e., better generalization capabilities of the trained model
Dependence on $K$. In distributed computing, scaling up typically results in proportional speed improvements. In DDL, however, scalability is less predictable due to the nuanced interplay of computation and communication costs with convergence, complicating the expected linear speedup. This unpredictability is starkly illustrated with VGG16* on the MNIST dataset across all data heterogeneity settings and all strategies. Specifically, Figure 6 demonstrates (at its upper part) that increasing the number of workers does not decrease computation – except for FedAdam which begins with significantly high computation – but rather exacerbates communication (shown at the upper part of the figure). These findings are troubling, as they reveal scaling up only hampers training speed and wastes resources. However, for more complex learning tasks like training DenseNet-121 and DenseNet-201 on CIFAR-10 (upper part of Figures 78), the expected behavior starts to emerge. Specifically, for DenseNet-121, scaling up ($K$ increase) leads to a decrease in computation cost for all strategies. Communication cost, however, increases with $K$ for all methods except Synchronous, which maintains constant communication irrespective of worker count, but at the expense of orders of magnitude higher communication overhead. Notably, while our findings might, in some cases, suggest potential speed benefits of not scaling up (smaller $K$), DDL is increasingly conducted within federated settings, where there is no other choice but to utilize the high number of workers.
FDA: Dependence on $\Theta$. The variance threshold $\Theta$ can be seen as a lever in balancing communication and computation; essentially, it calibrates the trade-off between these two costs. A higher $\Theta$ allows for greater model divergence before synchronization, reducing communication at the cost of potentially increased computation to achieve convergence. This impact of $\Theta$ is consistently observed across all two FDA strategies, learning tasks, and data heterogeneity settings (Figures 67 and 8). Interestingly, for more complex models like DenseNet121 and DenseNet201 on CIFAR-10, increasing the variance threshold ($\Theta$) does not lead to a significant rise in computation cost, as illustrated in Figures 7 and 8. It suggests that the FDA methods, by strategically timing synchronizations (monitoring the variance), substantially reduce the number of necessary synchronizations without a proportional increase in computation for the same model performance; this is particularly promising for complex DDL tasks.
Figure 6. VGG16* on MNIST — Accuracy Target: $0.994$
Figure 7. DenseNet121 on CIFAR-10 — Accuracy Target: $0.8$
Figure 8. DenseNet201 on CIFAR-10 — Accuracy Target: $0.78$
FDA: Choice of $\Theta$. The experimental results suggest that selecting any $\Theta$ within a specific order of magnitude (e.g., between $10^{2}$ and $10^{3}$ for DenseNet201) ensures convergence, as demonstrated in Figures 67, and 8. Therefore, identifying this range becomes crucial. To this end, we conducted extensive exploratory testing to estimate the $\Theta$ ranges for each learning task which are predominantly influenced by the number of parameters $d$ of the DNN. Within this context, $\Theta$ values outside the desirable range exhibit notable effects: below this range, the training process mimics Synchronous or Local-SGD approaches with small $\tau$, while exceeding it leads to non-convergence. Subsequently, having identified the optimal ranges for $\Theta$, we selected diverse values within them for our experimental evaluation (Table 1), thereby investigating different computation and communication trade-offs. For instance, in the ARIS-HPC environment with an InfiniBand connection (up to 56 Gb/s), experiments show that training wall-time (the total time required for the computation and the communication of the DDL) is predominantly influenced by computation cost, rendering communication concerns negligible. In such contexts, lower $\Theta$ values are favored due to their computational efficiency. On the contrary, in FL settings, where communication typically poses the greater challenge, opting for higher $\Theta$ values proves advantageous; reduction in communication achieved with higher $\Theta$ values will translate in a large reduction in total wall-time.
To assist researchers in selecting an optimal variance threshold, Figure 9 presents empirical estimations for $\Theta$ across three distinct learning settings: FL (assuming a common channel of $0.5$Gbps), Balanced (communication-computation equilibrium), and HPC.
$\Theta_{FL}$ $=4.91\cdot 10^{-5}\cdot d$
$\Theta_{B}$ $=3.89\cdot 10^{-5}\cdot d$
$\Theta_{HPC}$ $=2.74\cdot 10^{-5}\cdot d$
Figure 9. Empirical Estimation of the Variance Threshold
FDA: Linear vs. Sketch. In our main body of experiments, across most learning tasks and data heterogeneity settings, the two proposed FDA methods exhibit comparable performance, as illustrated in Figures 123, and  4. This suggests that the precision of the variance approximation is not critical; occasional ”unnecessary” synchronizations do not significantly impact overall performance. However, in all experiments within the more intricate transfer learning scenario, LinearFDA requires approximately $1.5$ times more communication than SketchFDA to fine-tune the deep ConvNeXtLarge model to equivalent performance levels (Figure 10).
In light of these findings, we conclude the following: for straightforward and less demanding tasks, LinearFDA is the recommended option due to its simplicity and lower complexity per local state computation. On the other hand, for intricate learning tasks and deeper models, SketchFDA becomes the preferred choice, particularly if communication-efficiency is paramount.
Figure 10. ConvNeXtLarge on CIFAR-100 (transfer learning from ImageNet) — Deployment of FDA during the fine-tuning stage with Accuracy Target of $0.76$
5. Conclusions
In this paper, we introduced Federated Dynamic Averaging (FDA), an innovative, adaptive and communication-efficient algorithm for distributed deep learning. Essentially, FDA makes informed, dynamic decisions on when to terminate training rounds based on approximations of the model variance. Through extensive experiments across diverse datasets and learning tasks, we demonstrated that FDA significantly reduces communication overhead (often by orders of magnitude) without a corresponding increase in computation or compromise in model performance — contrary to the typical trade-offs encountered in the literature. Furthermore, we showed that FDA is robust to data heterogeneity and inherently mitigates overfitting. Our results push the limits of modern communication-efficient distributed deep learning, paving the way for scalable, dynamic, and broadly applicable strategies.